A band teacher is taking some of her students to play at the town fair. If she is taking 1 trumpeter, 1 saxophonist, 1 drummer, 3 clarinetists, and 2 flutists, how many ways can she arrange them?
Answer by ikleyn(52879)

.
A band teacher is taking some of her students to play at the town fair. If she is taking 1 trumpeter, 1 saxophonist, 1 drummer,
3 clarinetists, and 2 flutists, how many ways can she arrange them?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

1 + 1 + 1 + 3 + 2 = 8 persons in total.

Answer. In 8! = 1*2*3*4*5*6*7*8 = 40320 ways.


This question is about the number of permutations.

On permutations, see the lesson Introduction to Permutations in this site.